OpenStudy (anonymous):

Find the sum of a finite arithmetic sequence from n = 1 to n = 15, using the expression 2n + 5. 315 215 32 35

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

Please help! Ill give medals and become fan

OpenStudy (anonymous):

what is the first term, when \(n=1\)?

OpenStudy (anonymous):

I dont know

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

replace \(n\) by \(1\) in \(2n+5\)

OpenStudy (anonymous):

so its 7?

OpenStudy (anonymous):

yes now hold that number

OpenStudy (anonymous):

what is \(2n+5\) if \(n=15\)?

OpenStudy (anonymous):

35?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Is that the final answer?

OpenStudy (anonymous):

formula you are going to use is \[\frac{n}{2}(a_1+a_n)\] you have \(a_1=7,a_{15}=35\)

OpenStudy (anonymous):

what is\(7+35\)?

OpenStudy (anonymous):

Its 42

OpenStudy (anonymous):

ok so we are at \[\frac{15}{2}(42)\]

OpenStudy (anonymous):

divide 42 by 2, then multiply the result by 15

OpenStudy (anonymous):

Its 315